# From Loops to Oops: Fallback Behaviors of Language Models Under Uncertainty

### v1

@@ 5 Factors Influencing the Fallback Behaviors of an LLM
Takeaway: While LLMs have some internal capability to avoid hallucinations, this fallback behavior is inherent to their generation scheme and is likely unavoidable with current decoding methods. Mitigating degenerate text through random sampling can come at the cost of more hallucinations.
@@ 5.1 Effect of sampling methods
Figure 6: Increasing the sampling temperature shifts fallback behavior. An average of 5 completions from Pythia-12B on TriviaFacts for varying temperatures. Green line marks the number of ground truth answers.
So far in our experiments, we mostly used greedy decoding to obtain responses from the models. However, in real-world applications, decoding often includes sampling from the model’s output distribution and including repetition/frequency penalties (e.g., (Holtzman et al., 2020; Keskar et al., 2019; Kumar et al., 2022)). We analyze the effect of temperature sampling on fallback behaviors by repeating our experiment on TriviaFacts, while generating five different sequences per model for each input, with increasing levels of randomness (using larger values of temperature). Figure 6 shows the results for Pythia-12B, revealing a surprising finding – while higher levels of randomness mitigate some repetitiveness and help models to break out of endless loops in the generation, [^5] This is expected as even when copying previous tokens with high probability such as $0.99$, continuing the generation long enough will inevitably lead to generating a token that is not part of the loop, thus steering the text to a new direction. it also causes a shift towards hallucinations. Moreover, we observe that introducing randomness reduces the number of correct facts, which could be attributed to the random skipping of correct facts associated with low confidence in the model.
@@ 5.2 Models “volunteer” hallucinations
In many previous experiments, we deliberately placed the model in high-uncertainty scenarios, prompting it to produce additional facts when none existed. To lift this restriction, we modify the base prompt in the TriviaFacts data by removing the requirement for a specific number of answers (25). For example, the modification to the example in the first row of Table 1 would result in the prompt ‘‘The following colors are in the Olympic rings:’’, allowing the model to stop generating additional answers once it exhausted its knowledge. However, the same trends continue to emerge (Figure 24), showcasing that even in natural situations without synthetically induced uncertainty, LLMs still fall back on the same behaviors and fail to use more appropriate “exit strategies”, such as topic changes or explicit abstaining.
@@ 5.3 Can fallbacks be prevented by prompting?
We modify the prompts as follows: For instruction-tuned models, we add the following prefix to each prompt ‘‘Complete the following list with facts you are sure of, and stop when you cannot recall additional facts’’. For pretrained models, we add a prefix with three demonstrations for in-context learning, each consists of an easy list question and its corresponding answers followed by a topic-change to a new list (Figure 25 gives the full prefix). While for Pythia and OLMo we observe a minor increase in abstaining behavior at the expense of hallucinations, the overall trend remains the same, with hallucinated facts emerging abundantly (Figure 26). In comparison, both Llama 2 and more pronouncedly Llama 3 manage to abstain by changing topics often, but the portion of hallucinated facts still increases with model size. We conclude that fallback behaviors are inherent to current pretrained LLMs and emerge “unintentionally” and unavoidably under uncertainty.
@@ 6.1 Emergence of fallbacks during generation
Figure 7: Models shift fallback behavior during a single completion. Order of fallbacks per generation for a Pythia-12B model—each row represents a specific prompt with the 25 produced facts. Green marks correct answers, orange hallucinations and blue repeated facts. Questions are sorted by the number of consecutive repetitions.
We look at the facts generated by the model in response to a given query as an ordered list of labels (i.e. correct, hallucination, repetition). For example, Figure 7 depicts in each row the 25 labels of the facts produced by Pythia-12B for each of the 95 samples in TriviaFacts. Surprisingly, we observe that almost without exception, the model first generates all the correct facts it can (green), then shifts to producing hallucinations (orange), and finally shifts further to repeating previous facts (blue). This is especially noteworthy as we see that the fallback behavior is not determined by the question, and the ordering between fallbacks follows the same ordering as when increasing model strength. This trend holds regardless of the model, the dataset, or the decoding method (see Figures 32, 33, 34, 35, 36, 37, 38 and 39 for further results).
To quantify this phenomenon, we denote a label by 1 for repetition, 2 for hallucination, and 3 for correct, and define
@@
Where $f_{1},\dots,f_{n}$ is an ordered list of fact-labels ($f_{i}\in\{1,2,3\}$), and $\mathbbm{1}$ is the indicator function. For each pair of dataset and model, we measure the ShiftScore for each generation. We then run a Mann-Whitney U-test (McKnight and Najab, 2010) to compare the results to the expected ShiftScore if the list of facts was produced in a random order (more details in Section D.2). We find that for all tested pairs, the p-value is $<10^{-8}$ (see Table 2 for full analysis), and the U-statistic is always positive. We thus conclude that the ordering is not random, and follows the aforementioned hierarchy of fallback behaviors.
When investigating the fallback behaviors in a single generation, one may expect to observe occurrences of verbatim recollection of training samples, as suggested by Nasr et al. (2023a). To address this, we analyzed the generations from our experiments both qualitatively and quantitatively, and found no evidence of such behavior. We therefore hypothesize that verbatim generation of training samples stems from other factors than uncertainty regarding factual information, and is more prevalent in the presence of out-of-distribution inputs. For further details, see Appendix C.
@@ 6.2 Fallback shifts is a continuous process
Figure 8: Model output gradually degenerates with increased generation length. When generating biographies for rare entities, model output degenerate as completion length increases, compared to the human baseline (Wikipedia entries for the same entities). Stronger models’ output deteriorates more slowly.
When analyzing generations as a discrete list of facts, we also find discrete shifts in fallback behaviors as the model’s uncertainty grows. While examining strict repetition of facts is of particular interest, one may also consider a softer measure of degenerate text generation. For example, producing a list of URLs such as https://page.domain/xyz, https://page.domain/xyq, https://page.domain/wyz is not strictly a repetition nor necessarily a hallucination, but it is clear that some form of conditional repetitiveness is occurring. Similarly, producing the same entities in random permutations is clearly undesired behavior and is not strictly a repeated sequence. For example, in Figure 12 we can see that the model starts repeating previous facts such as the shows Elsa Pataky appeared in, though with a slightly different phrasing, before it begins repeating previous sequences verbatim.
To measure such phenomena, we introduce a DiversityScore. Given a sequence of predicted tokens $x=t_{1},\dots,t_{n}$ from the model’s vocabulary $\mathcal{V}$, we define $\texttt{DiversityScore}(x)\coloneqq\frac{1}{n}\sum_{v\in\mathcal{V}}{\mathbbm{1}_{v\in x}}$, i.e. the number of unique tokens in the sequence divided by the sequence length. When $n\ll|\mathcal{V}|$ we expect $\texttt{DiversityScore}\approx 1$ while a repetitive sequence will have a DiversityScore that diminishes rapidly towards 0.
When experimenting with naturally occurring (human-generated) sequences, such as Wikipedia paragraphs, for sequences up to $512$ tokens, the DiversityScore is almost always higher than $0.6$ and often ranges around $0.7-0.8$ even for longer sequences (see Figure 8). Specifically, we use the subset of BioGeneration entities that are very frequent according to Min et al. (2023b), as we expect the least amount of uncertainty to appear and the most full and diverse generations, as discussed in Section 4.4. We compare the Wikipedia paragraphs for these entities to models’ outputs when generating biographies for them. For all models’ outputs, the DiversityScore diminishes much faster, with weaker models crossing the $0.5$ threshold at around 150 tokens (i.e., after only 150 tokens, every other token is repeating a previous one). We include the same analysis on generations from the stronger model Llama 3 8B (AI, 2024) for comparison. Indeed, while it produces slightly more diverse outputs, diversity diminishes towards zero as the sequence length grows.
@@ 7 Related Work
Holtzman et al. (2020) first attributed the tendency of LMs to generate highly repetitive and degenerate text to greedy sampling and suggested nucleus sampling as a possible mitigation. Follow-up work demonstrated the role of uncertainty in generating degenerate text and proposes various solutions using random, controlled, or constrained generation techniques Keskar et al. (2019); Kumar et al. (2022); Zhang et al. (2022); Finlayson et al. (2023); Su et al. (2022). In a parallel approach, Olsson et al. (2022) focused on understanding the intrinsic mechanisms that cause models to copy previous inputs.
@@
As random sampling techniques became ubiquitous and LLMs grew in size and capability, the main focus shifted toward understanding the causes and proposing solutions to the generation of non-factual text, commonly referred to as hallucinations. Recent work has focused on understanding how and why hallucinations emerge during generation Rashkin et al. (2023); Zhang et al. (2023); Adlakha et al. (2023); Kim et al. (2024), reducing hallucinations (e.g., by retrieval-augmentation or prompting) Roller et al. (2021); Shuster et al. (2021); Dhuliawala et al. (2023), and detecting hallucinations Zhou et al. (2021); Liu et al. (2022); Honovich et al. (2022); Min et al. (2023b); Mishra et al. (2024); Gottesman and Geva (2024). For a survey on hallucinations in natural language generation, see Ji et al. (2022). In concurrent work, Denison et al. (2024, inter alia) demonstrated how models generalize during training and learn to shift from simple dishonest strategies such as sycophancy to generating false facts with premeditation. Similarly, Band et al. (2024) suggest that the problem of hallucinating rather than abstaining is an issue with the calibration of the model, as was also hypothesized by OpenAI et al. (2024).
@@
The phenomenon where LMs abstain or hallucinate when uncertain is well-known and has been reported in previous work Xiao and Wang (2021); Lin et al. (2022); Snyder et al. (2023); Baan et al. (2023); Kang et al. (2024). Recent studies have examined whether the generation probability of LMs is calibrated Kadavath et al. (2022) and whether they can express their uncertainty in natural language Li et al. (2023); Yona et al. (2024). Unlike previous research, our focus is on understanding the varying factors impacting the different fallback behaviors of LMs under uncertainty.
8 Conclusion
This work links the notorious unwanted behaviors of LLMs, such as degenerate and repetitive text and hallucinations, showing that they are all fallback behaviors for models when faced with uncertainty. We provide abundant evidence that these behaviors emerge with a clear ordering between their appearances, either when comparing similar models of different strength, different decoding strategies and even in a single generation. Our work suggests that these fallback behaviors are inherent to current LLMs and that the ubiquitous methods to alleviate such issues may simply replace one fallback by another. Moreover, increasing model strength by training longer or adding more parameters enhances overall performance and shifts its fallbacks towards more complex range. However, as generation length grows, even the strongest models will resort to hallucinations and may eventually produce degenerate text.
@@ 8 Conclusion
Second, while we study the effect of random sampling, it becomes more common to apply even stricter modifiers to the language model next-word prediction distribution by using methods such as nucleus sampling, top-k decoding, repetition penalty and more (Holtzman et al., 2020; Keskar et al., 2019; Finlayson et al., 2023). We leave to future work to examine such decoding strategies on fallback behaviors.
@@ Acknowledgments
Appendix A Datasets
@@ A.3 Qampari
A.4 FakeQampari
We take the questions from Qampari (see Section A.3) and employ ChatGPT to replace the subject entity in each question with an alternative entity name which sounds plausible but does not exist (for example, choosing a Japanese name for an anime creator). We manually verified the generated questions, while checking that all new subjects feel plausible in the context and that they refer to entities that do not exist.
Appendix B Parsing Answers

### v2

@@ 5 Factors Influencing the Fallback Behaviors of an LLM
Takeaway: While LLMs have some internal capability to avoid hallucinations, this fallback behavior is inherent to their generation scheme and is likely unavoidable with current decoding methods. Mitigating degenerate text through random sampling often comes at the cost of more hallucinations.
@@ 5.1 Effect of sampling methods
Figure 5: Larger models hallucinate more when generating biographies of rare entities. Larger Pythia models produce more atomic facts and more hallucinations on Bio.
Figure 6: Increasing the sampling temperature shifts fallback behavior. An average of 5 completions from Pythia-12B on TriviaFacts for varying temperatures.
So far we mostly used greedy decoding to obtain responses from the models. However, in real-world applications, decoding often includes sampling from the model’s output distribution and including repetition/frequency penalties (e.g., Holtzman et al., 2020; Keskar et al., 2019; Kumar et al., 2022). We analyze the effect of temperature sampling on fallback behaviors by repeating our experiment on TriviaFacts, while generating five sequences per model for each input with an increasing temperature. Figure 6 shows the results for Pythia-12B, surprisingly revealing that while a higher temperature mitigates some repetitiveness, it causes a shift towards hallucinations. Moreover, introducing randomness reduces the number of correct facts, [^7] For example, if the first token for a correct fact has a probability of 0.51, it would be selected in greedy decoding but may be skipped with random sampling. which could be attributed to the random skipping of correct facts that have low confidence in the model. When repeating this experiment with additional models (Figures 29, 30, 31, 32 and 33), as well as when using top-p (“nucleus”) sampling (Holtzman et al., 2020), the results produce an identical trend. This further demonstrates the strong relationship between the fallback behaviors.
@@ 5.2 Models “volunteer” hallucinations
In many previous experiments, we deliberately placed the model in high-uncertainty scenarios, prompting it to produce additional facts when none existed. To lift this restriction, we modify the base prompt in the TriviaFacts data by removing the requirement for a specific number of answers (25). For example, the modification to the example in the first row of Table 1 would result in the prompt ‘‘The following colors are in the Olympic rings:’’, allowing the model to stop generating additional answers once it exhausted its knowledge. The same trends persist: larger Pythia models reduce repetitions from $16.4$ to $9.7$ but increase hallucinations from $2.9$ to $6.2$ (Figure 24). This shows that even in natural scenarios without synthetic uncertainty, LLMs still rely on the same behaviors, failing to use better “exit strategies” like topic changes or abstention.
@@ 5.3 Can fallbacks be prevented by prompting?
We modify the prompts as follows: For instruction-tuned models, we add the following prefix to each prompt ‘‘Complete the following list with facts you are sure of, and stop when you cannot recall additional facts’’. For pretrained models, we add a prefix with three demonstrations for in-context learning, each consists of an easy list question and its corresponding answers followed by a topic-change to a new list (Figure 16 gives the full prefix).
The results for base models on TriviaFacts and Qampari are presented in Figures 25 and 26, respectively. While for Pythia and OLMo we observe a minor increase in abstaining behavior at the expense of hallucinations, the overall trend remains the same, with hallucinated facts emerging abundantly (over $7$ on average). In comparison, both Llama 2 and more pronouncedly Llama 3 manage to abstain by changing topics often (up to $20$ facts on average), but the portion of hallucinated facts still increases with model size. Interestingly, when instructed to generate only facts that the model is certain about, nearly all instruction-tuned models generated even more hallucinations compared to their non-prompted counterparts, presumably indicating that they are “sure” of the facts they produce (Figures 27 and 28). Despite reacting differently to this modified instruction, all instruction-tuned models continued to exhibit similar fallback tendencies compared to when no explicit abstaining instruction was given. We conclude that fallback behaviors are inherent to current pretrained LLMs and emerge “unintentionally” and unavoidably under uncertainty.
@@ 6 Fallback Behaviors in One Generation
Figure 7: Models shift fallback behavior during generation. Order of fallbacks per generation for a Pythia-12B—each row shows a prompt with 25 facts from Triv. Green marks correct answers, orange hallucinations, blue repetitions, and red errors. Rows are sorted by consecutive repetition.
Figure 8: Model output gradually degenerates with increased generation length. When generating biographies for rare entities, output degenerates as completion length increases compared to the human baseline (respective Wikipedia entries). Stronger models deteriorate more slowly.
@@ 6.1 Emergence of fallbacks during generation
We view the facts generated by the model for a query as an ordered list of labels (correct, hallucination, repetition). For example, each row in Figure 8 shows the 25 labels of facts produced by Pythia-12B for each of the 95 samples in TriviaFacts. Surprisingly, the model almost always first generates correct facts (green), then shifts to hallucinations (orange), and finally to repeating facts (blue). Notably, this fallback behavior isn’t determined by the question and follows the same order as when increasing model strength. This trend holds regardless of the model, the dataset, or the decoding method (see Figures 40, 40, 42, 42, 44, 44, 46 and 46 for further results).
To quantify this phenomenon, we denote a label of 1 for repetition, 2 for hallucination, and 3 for correct, and define
@@
Where $f_{1},\dots,f_{n}$ is an ordered list of fact-labels ($f_{i}\in\{1,2,3\}$), and $\mathbbm{1}$ is the indicator function. For each pair of dataset and model, we measure the ShiftScore for each generation. We then run a Mann-Whitney U-test (McKnight & Najab, 2010) to compare the results to the expected ShiftScore if the list of facts was produced in a random order (more details in Section E.2). We find that for all tested pairs, the p-value is $<10^{-8}$ (see Table 2 for full analysis), and the U-statistic is always positive. We thus conclude that the ordering is not random, and follows the aforementioned hierarchy of fallback behaviors.
When investigating the fallback behaviors in a single generation, one may expect to observe occurrences of verbatim recollection of training samples, as suggested by Nasr et al. (2023a). To address this, we analyzed the generations from our experiments both qualitatively and quantitatively, and found no evidence of such behavior. We therefore hypothesize that verbatim generation of training samples stems from other factors than uncertainty regarding factual information, and is more prevalent in the presence of out-of-distribution inputs. For further details, see Appendix D.
@@ 6.2 Fallback shifts is a continuous process
So far, we have analyzed model generations as discrete lists of facts, and correspondingly observed discrete shifts in fallback behaviors. In this section, we consider a softer measure of degenerate text generation. For example, producing a list of URLs such as https://page.domain/xyz, https://page.domain/xyq, https://page.domain/wyz is not strictly a repetition nor necessarily a hallucination, but it is clear that some form of conditional repetitiveness is occurring. Figure 12 shows another example as the model starts repeating previous facts (such as the shows Elsa Pataky appeared in) with a slightly different phrasing, before it begins repeating previous sequences verbatim. We consider degenerate text as sequences with repetitive textual patterns and/or rephrasing of previously generated text. Common patterns include enumerated lists of partially repeated items, repetitions of names and sentences with minor perturbations and permutation of previously generated statements. At it’s peak, degenerate text is manifested by full verbatim repetition (Holtzman et al., 2020).
To measure such phenomena, we introduce a DiversityScore. Given a sequence of generated tokens $x=t_{1},\dots,t_{n}$ from the model’s vocabulary $\mathcal{V}$, we define $\texttt{DiversityScore}(x)\coloneqq\frac{1}{n}\sum_{v\in\mathcal{V}}{\mathbbm{1}_{v\in x}}$, i.e. the number of unique tokens in the sequence divided by the sequence length. When $n\ll|\mathcal{V}|$ we expect $\texttt{DiversityScore}\approx 1$ while a repetitive sequence will have a score that diminishes rapidly towards 0.
Experimenting with human-written texts of $\leq 512$ tokens from Wikipedia, we observe they have a typical DiversityScore of $0.7-0.8$ that almost never exceeds $0.6$ (see Figure 8). Specifically, we use the subset of frequent entities in BioGeneration Min et al. (2023b), for which we expect models to be the least uncertain and generate full and diverse outputs, as discussed in Section 4.4. We compare the Wikipedia paragraphs for these entities with the model-generated biographies. For all models’ outputs, the DiversityScore diminishes much faster, with weaker models crossing the $0.5$ threshold at around 150 tokens (i.e., after only 150 tokens, every other token appeared before). We observe a similar trend for the stronger model Llama 3 8B (Meta AI, 2024).
@@ 7 Related Work
Holtzman et al. (2020) first attributed the tendency of LMs to generate highly repetitive and degenerate text to greedy sampling and suggested nucleus sampling as a possible mitigation. Follow-up work demonstrated the role of uncertainty in generating degenerate text and proposes various solutions using random, controlled, or constrained generation techniques Keskar et al. (2019); Kumar et al. (2022); Zhang et al. (2022); Finlayson et al. (2023); Su et al. (2022); Li et al. (2023a). In a parallel approach, Olsson et al. (2022) focused on understanding the intrinsic mechanisms that cause models to copy previous inputs.
@@
As random sampling techniques became ubiquitous and LLMs grew in size and capability, the main focus shifted toward understanding the causes and proposing solutions to the generation of non-factual text, commonly referred to as hallucinations Ji et al. (2022); Huang et al. (2023). Recent work has focused on understanding how and why hallucinations emerge during generation Rashkin et al. (2023); Zhang et al. (2023); Adlakha et al. (2023); Kim et al. (2024), reducing hallucinations (e.g., by retrieval-augmentation or prompting) Roller et al. (2021); Shuster et al. (2021); Dhuliawala et al. (2023), and detecting hallucinations Zhou et al. (2021); Liu et al. (2022); Honovich et al. (2022); Min et al. (2023b); Mishra et al. (2024); Gottesman & Geva (2024); Hron et al. (2024). Recently, Denison et al. (2024) demonstrated how models generalize during training and learn to shift from simple dishonest strategies such as sycophancy to generating false facts with premeditation. Similarly, Band et al. (2024) suggest that hallucinating rather than abstaining is an issue with the calibration of the model, as was also hypothesized by OpenAI et al. (2024).
@@
The phenomenon where LMs abstain or hallucinate when uncertain is well-known Xiao & Wang (2021b); Lin et al. (2022); Snyder et al. (2023); Baan et al. (2023); Kang et al. (2024). Recent studies have examined if the generation probability of LLMs is calibrated Kadavath et al. (2022) and whether they can express their uncertainty in natural language Li et al. (2023b); Yona et al. (2024). A parallel line of work focuses on defining notions of uncertainty and designing method to quantify it in various setups (Xiao & Wang, 2021a; Hou et al., 2023; Lin et al., 2023; Liu et al., 2024b; Ling et al., 2024). Unlike previous research, we focus on understanding model behaviors under epistemic uncertainty, rather on quantifying such uncertainty.
Specifically, we focus on epistemic uncertainty that reflects a lack of knowledge, regardless of its root cause, such as training data limitations, model architecture, or optimization strategies. This approach aligns with the definitions provided by Hüllermeier & Waegeman (2019), where epistemic uncertainty is characterized as reducible through additional data or better modeling, and complements the broader uncertainty taxonomy discussed by Baan et al. (2023). Additionally, Abbasi-Yadkori et al. (2024) emphasize the importance of distinguishing epistemic uncertainty from aleatoric uncertainty, which arises from inherent randomness, particularly in complex outputs or out-of-distribution settings. While our study primarily investigates epistemic uncertainty in knowledge-intensive use-cases of LLMs, it would also be valuable for future work to explore aleatoric uncertainty and its implications for model behavior in cases of ambiguous or out-of-distribution inputs.
8 Conclusion and Discussion
This work links the notorious unwanted behaviors of LLMs, such as degenerate and repetitive text and hallucinations, showing that they are all fallback behaviors models exhibit under uncertainty. We provide abundant evidence that these behaviors emerge with a clear ordering between their appearances, when comparing similar models of different strength, different decoding strategies and even in a single generation. Our experiments suggest that these fallback behaviors are inherent to current LLMs and that existing methods to alleviate them may simply replace one fallback by another. Moreover, longer training and additional parameters enhance performance and shift model fallbacks towards more complex range. However, as generation length grows, even the strongest models will resort to hallucinations and may eventually produce degenerate text.
Our findings have practical implications for the deployment and scalable oversight of LLMs (Bowman et al., 2022). Specifically, understanding fallback behaviors under epistemic uncertainty can inspire practitioners with actionable insights into how models respond when faced with gaps in knowledge. This is particularly critical as model sizes increase, given the observed tendency of larger models to exhibit subtler fallback behaviors, such as hallucinations, which are more challenging to detect compared to simpler repetitions. Future research may explore mechanisms to steer models toward predictable fallback behaviors, leveraging the connections between behaviors we identify to develop practical interventions for reducing hallucinations in deployed systems (Mosbach et al., 2024).
@@ 8 Conclusion and Discussion
Second, while we study the effect of random sampling, it becomes more common to apply even stricter modifiers to the language model next-word prediction distribution by using methods such as top-k decoding, repetition penalty, frequency penalty and more (Keskar et al., 2019; Finlayson et al., 2023). We leave to future work to examine such decoding strategies on fallback behaviors.
@@ Acknowledgments
Appendix
Appendix A Reproducibility
We have made every effort to ensure the reproducibility of our work. All code, datasets, hyperparameters, and recipes required to reproduce our results and plots are available in our https://github.com/Mivg/fallbacks. This repository also contains all model generations and the complete set of plots, some of which are presented in the paper. None of the models or data used are proprietary; all experiments rely on open-source frameworks, including HuggingFace and PyTorch, and publicly available datasets. Detailed descriptions of the experimental setup and data processing steps can be found in the supplementary materials. Specifically, Section 3 outlines the experimental setup and the process for generating answers from various models. Appendix B describes the collection and processing procedures for each dataset used in the study. Additionally, Appendix C provides implementation details on the parsing process, including fact extraction and classification for each type of generated output.
Appendix B Datasets
@@ B.3 Qampari
B.4 FakeQampari
We take the questions from Qampari (see Section B.3) and employ ChatGPT to replace the subject entity in each question with an alternative entity name which sounds plausible but does not exist (for example, choosing a Japanese name for an anime creator). We manually verified the generated questions, while checking that all new subjects feel plausible in the context and that they refer to entities that do not exist.
Appendix C Parsing Answers